OpenStudy (anonymous):

integral of (sqrt)1+sin(theta)

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\sqrt{1+\sin(\theta)}d \theta = -\cos/\sqrt{1-\sin(\theta)}\]

OpenStudy (anonymous):

this is what they get

OpenStudy (anonymous):

have i stumped everone haha

OpenStudy (dumbcow):

i believe this cannot be integrated. at least as an indefinite integral

OpenStudy (dumbcow):

the answer above is wrong because its derivative does not equal sqrt(1+sin)

OpenStudy (anonymous):

I have a very good thing about this one \[I = \int_0^{2 \pi} \sqrt{1 +sin\theta}d \theta\] \[I = \int_0^{2\pi} \sqrt{1 + sin{2\pi - \theta}}d\theta\] \[I^2 = \int_0^{2\pi} \sqrt{1 - sin^2\theta}d\theta\] \[I^2 = \int_0^{2\pi}cos\theta d\theta\] \[I^2 = sin\theta|_0^{2\pi}\] \[I = 0\]

OpenStudy (dumbcow):

ahh good answer how did you get sin^2 ?? sin(2pi -theta) = -sin(theta)

OpenStudy (anonymous):

right ... but wolfram doesn't agree with me 4sqrt2 is the wolfram answer

OpenStudy (dumbcow):

oh you squared I

OpenStudy (anonymous):

yeah multiplied two equations maybe i have to find the integral I not I that means sqrtcos{theta}..

OpenStudy (anonymous):

wolfram confuses me more ah it says complex answer to sqrtcosx form 0 to 2pi

OpenStudy (dumbcow):

yeah its 4sqrt(2), i just graphed it

OpenStudy (anonymous):

wolfram one